The concentration of X²⁻ in a saturated solution of A₂X₃ is 9.62 x 10⁻¹¹ mg/L. What is the Ksp of
A₂X₃ if the MM of X is 32.06 g/mol?
1
Expert's answer
2020-12-30T04:02:43-0500

A2X3 = 2 A3+ + 3 X2-

Ksp = [A3+]2 × [X2-]3

since concentration of both the ions are same.

[X2-] = (9.62×10-14 g/L)/(32.06 g/mol)

[X2-] = 3 × 10-15 M


so Ksp = (3×10-15)5 = 2.43 × 10-73